If: Y = C + I where C = 120 + 0.8Y, I = 200 + 0.05i and Y = 1400,
Q1. Compute the equilibrium national income, consumption and investments
Q2. What is the size of government expenditure multiplier and savings ?
1
Expert's answer
2015-03-19T11:32:01-0400
Q1.
Y = C + I = 120 + 0.8Y + 200 + 0.05i = 320 + 0.8Y+ 0.05i
0.2Y = 320+ 0.05i
280 – 320 = 0.05i
i = -800
I = 200 + 0.05i = 200 – 40 = 360
C = 120 + 0.8Y = 120 + 1120 = 1240
Y = C + I = 1400

Q2.
government expenditure multiplier = 1 / (1-MPC)
MPC is the marginal propensity to consume
government expenditure multiplier = 1 / (1-0.8) = 5
We have the consumption function in the form C = a + MPC (Y) that
autonomous consumption is a. The savings function is S = -a + (1 –
MPS) (Y)
So, the saving function is:
S = -120 + 0.2 Y
S = -120 + 280 = 160